OpenStudy (anonymous):

Factor the expression completely. (a2 + 1)2 − 43(a2 + 1) + 442

OpenStudy (anonymous):

a^4+2a^2+1-43a^2-43+442 =a^4-41a^2+400 Let x=a^2 x^2-41x+400 make it equal 0 and solve it 41+-sqrt(1681-1600)/2 x1=25 x2=16 So a^2=25 a=+-5 or a^2=16 a=+-4 So roots of a^4-41a+400 are -5,-4,4,5 so it factors to (a+5)(a+4)(a-4)(a-5)